\begin{thebibliography}{Kaw92}

\bibitem[BS06]{Birkar-Shokurov}
C. Birkar,  V. V. Shokurov.  Mld's vs thresholds and flips.
E-print math.AG/0609539.

\bibitem[Cor95]{Corti1995}
A. Corti.
\newblock {Factoring birational maps of threefolds after {S}arkisov}.
\newblock {\em J. Algebr. Geom.}, 4(2):223--254, 1995.

\bibitem[Kaw92]{Kawamata-1992-discr}
Y. Kawamata.
\newblock The minimal discrepancy coefficients of terminal singularities in
  dimension three. {A}ppendix to {V}.{V}. {S}hokurov's paper {"}3-fold log
  flips{"}.
\newblock {\em Russ. Acad. Sci., Izv., Math.}, 40(1):95--202, 1992.

\bibitem[KM98]{Kollar-Mori-19988}
J. Koll{\'a}r and S. Mori.
\newblock {\em Birational geometry of algebraic varieties}, volume 134 of {\em
  Cambridge Tracts in Mathematics}.
\newblock Cambridge University Press, Cambridge, 1998.
\newblock With the collaboration of C. H. Clemens and A. Corti, Translated from
  the 1998 Japanese original.

\bibitem[Kol92]{Utah}
J.~Koll{\'a}r, editor.
\newblock {\em Flips and abundance for algebraic threefolds}.
\newblock Soci\'et\'e Math\'ematique de France, Paris, 1992.
\newblock Papers from the Second Summer Seminar on Algebraic Geometry held at
  the University of Utah, Salt Lake City, Utah, August 1991, Ast\'erisque No.
  211 (1992).

\bibitem[Kol94]{Kollar1994}
J. Koll{\'a}r.
\newblock Log surfaces of general type; some conjectures.
\newblock In {\em Classification of algebraic varieties (L'Aquila, 1992)},
  volume 162 of {\em Contemp. Math.}, pages 261--275. Amer. Math. Soc.,
  Providence, RI, 1994.

\bibitem[Kol97]{Kollar-1995-pairs}
J. Koll{\'a}r.
\newblock Singularities of pairs.
\newblock In {\em Algebraic geometry---Santa Cruz 1995}, volume~62 of {\em
  Proc. Sympos. Pure Math.}, pages 221--287. Amer. Math. Soc., Providence, RI,
  1997.

\bibitem[Kol08]{kollar-2008}
J. Koll\'ar.
\newblock Which powers of holomorphic functions are integrable? Preprint,
arXiv.org:0805.0756, 2008.

\bibitem[Mar96]{Markushevich-1996-discr}
D. Markushevich.
\newblock Minimal discrepancy for a terminal c{DV} singularity is {$1$}.
\newblock {\em J. Math. Sci. Univ. Tokyo}, 3(2):445--456, 1996.

\bibitem[MP04]{Prokhorov-McKernan-2004}
J. McKernan and Y. Prokhorov.
\newblock Threefold thresholds.
\newblock {\em Manuscripta Math.}, 114(3):281--304, 2004.

\bibitem[Rei80]{Reid-1980can}
M. Reid.
\newblock Canonical {$3$}-folds.
\newblock In {\em Journ\'ees de G\'eometrie Alg\'ebrique d'Angers, Juillet
  1979/Algebraic Geometry, Angers, 1979}, pages 273--310. Sijthoff \&
  Noordhoff, Alphen aan den Rijn, 1980.

\bibitem[Rei87]{Reid-YPG1987}
M. Reid.
\newblock Young person's guide to canonical singularities.
\newblock In {\em Algebraic geometry, Bowdoin, 1985 (Brunswick, Maine, 1985)},
  volume~46 of {\em Proc. Sympos. Pure Math.}, pages 345--414. Amer. Math.
  Soc., Providence, RI, 1987.

\bibitem[Sho88]{Shokurov1988p}
V.~V. Shokurov.
\newblock Problems about {F}ano varieties.
\newblock In {\em Birational Geometry of Algebraic Varieties, Open Problems},
  pages 30--32, Katata, 1988.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb,mathrsfs, stmaryrd}%,mathcomp,amsfonts, array}

\makeatletter \@addtoreset{equation}{section} \makeatother

\newcommand{\down}[1]{\left\lfloor #1\right\rfloor}
\newcommand{\muu}{{\boldsymbol{\mu}}}
\newcommand{\alphaa}{{\boldsymbol{\alpha}}}

\newcommand{\CC}{{\mathbb C}}
\newcommand{\QQ}{{\mathbb Q}}
\newcommand{\OOO}{{\mathscr{O}}} 

\newcommand{\ord}{\operatorname{ord}}

\newcommand{\ct}{\operatorname{ct}}

\newcommand{\Tc}{{\mathcal T}^{\mathrm{can}}}

\newcommand{\wt}{\operatorname{wt}}

\newtheorem{theorem}[equation]{Theorem}

\newtheorem{lemma}[equation]{Lemma}

\newtheorem{conjecture}[equation]{Conjecture}

\theoremstyle{definition}
\newtheorem{definition}[equation]{Definition}

\newtheorem{example}[equation]{Example}
\newtheorem{case}[equation]{}

\author{Yuri Prokhorov}
\thanks{
The author was
partially supported by 
RFBR, Nos. \ 08-01-00395-a and 06-01-72017-MHTI-a.
}
\address{
Yuri Prokhorov, Department of Higher Algebra, 
Faculty of Mathematics and Mechanics, 
Moscow State Lomonosov University, Vorobievy Gory, Moscow, 
119 899, RUSSIA}
\email{prokhoro@mech.math.msu.su} 
\title{Gap conjecture for $3$-dimensional canonical 
thresholds}
\subjclass{14B05, 14J17, 14E30, 32S25}
\date{}

\begin{document}
\maketitle
\begin{abstract}
We prove that the interval $(5/6,\, 1)$ contains no $3$-dimensional canonical 
thresholds.
\end{abstract}

\section{Introduction}

Let $(X\ni P)$ be a three-dimensional 
canonical singularity 
and let $S\subset X$ be a $\QQ$-Cartier divisor.
The \emph{canonical threshold} of the pair $(X,S)$ is 
\[
\ct(X,S):=\sup \{ c \mid \text{the pair $(X,cS)$ is canonical}\}.
\]
It is easy to see that $\ct(X,S)$ is rational and non-negative. 
Moreover, if $S$ is effective and integral, then $\ct(X,S)\in [0,\, 1]$.
Define the subset $\Tc_n\subset [0,\, 1]$ as follows
\[
\Tc_n:=\{\ct(X,S) \mid \text{$\dim X=n$, $S$ is 
integral and effective} 
\}.
\]

The following conjecture is an analog of 
corresponding conjectures for log canonical 
thresholds and minimal discrepancies, see \cite{Shokurov1988p},
\cite{Utah}, \cite{Kollar-1995-pairs}, \cite{Prokhorov-McKernan-2004}, \cite{kollar-2008}.

\begin{conjecture}
\label{conj-acc}
The set $\Tc_n$ satisfies the ascending chain condition.
\end{conjecture}
The conjecture is interesting for applications to 
birational geometry, see, e.g., \cite{Corti1995}.
It was shown in \cite{Birkar-Shokurov} that much more general form of \ref{conj-acc}
follows from ACC for minimal log discrepancies 
and weak Borisov-Alexeev conjecture.
The important particular case of \ref{conj-acc} is the following
\begin{conjecture}[cf. {\cite{kollar-2008}}]
\label{conj-gap}
$\epsilon_n^{\mathrm{can}}:=1-\sup (\Tc_n\setminus \{1\})>0$.
\end{conjecture}
\noindent
The aim of this note is to prove Conjecture \ref{conj-gap} for $n=3$ in a 
precise form:

\begin{theorem}
\label{main-1}
$\epsilon^{\mathrm{can}}_3=1/6$.
\end{theorem}
An analog of this theorem for log canonical thresholds was
proved by J. Koll\'ar \cite{Kollar1994}: $\epsilon^{\mathrm{lc}}_3=1/42$.

Note that replacing $(X\ni P)$ with its terminal $\QQ$-factorial
modification we may assume that $(X\ni P)$ is terminal.
Thus the following is a stronger form of Theorem \ref{main-1}:

\begin{theorem}
\label{main-2}
Let $(X\ni P)$ be a three-dimensional terminal singularity 
and let $S\subset X$ be an \textup(integral\textup) 
effective Weil $\QQ$-Cartier divisor
such that the pair $(X,S)$ is not canonical.
Then $\ct(X,S)\le 5/6$. 
Moreover, if $(X\ni P)$ is singular, then $\ct(X,S)\le 4/5$. 
\end{theorem}
The proof is rather standard.
We use the classification of terminal singularities and 
weighted blowups techniques, cf. \cite{Kawamata-1992-discr}, \cite{Kollar1994},
\cite{Markushevich-1996-discr}.

\section{Preliminaries}

\begin{case} \textbf{Notation.}
For a polynomial $\phi$, 
$\ord_0 \phi$ denotes the order of vanishing of 
$\phi$ at $0$ and
$\phi_d$ is the homogeneous 
component of degree $d$.

Throughout this paper we let $(X\ni P)$ be the germ of 
a three-dimensional terminal singularity and let 
$S\subset X$ be an 
effective Weil $\QQ$-Cartier divisor
such that the pair $(X,S)$ is not canonical.
Put $c:=\ct(X,S) >0$. Since $(X,S)$ is not canonical, $c<1$.

We work over the complex number field $\CC$.
\end{case}

\begin{lemma}
In the above notation the singularity $(S\ni P)$ is not Du Val.
\end{lemma}

\begin{proof}
This is well-known, see e.g. \cite[Th. 2.6]{Reid-1980can}.
\end{proof}

\begin{case}
We use the techniques of weighted blowups.
For definitions and basic properties we refer, for example, 
to \cite{Markushevich-1996-discr}, \cite{Reid-YPG1987}.
By fixing coordinates $x_1,\dots,x_n$
we regard the affine space $\CC^n$ as a toric variety.
Let $\alphaa=(\alpha_1,\dots,\alpha_n)$ be a weight
(a primitive lattice vector in the positive octant)
and let $\sigma_{\alpha}\colon \CC^n_{\alphaa}\to \CC^n$
be the weighted blowup with weight $\alphaa$ ($\alphaa$-blowup).
The exceptional divisor $E_{\alphaa}$ is irreducible and 
determines a discrete valuation $v_{\alphaa}$ of the function field 
$\CC(\CC^n)$ such that $v_{\alphaa}(x_i)=\alpha_i$.
\end{case}

\begin{case}
Now let $X\subset \CC^n$ be a hypersurface given by 
the equation $\phi=0$ and let $X_{\alphaa}\subset \CC^n_{\alphaa}$
be its proper transform.
Fix an irreducible component $G$ of $E_{\alphaa}\cap X_{\alphaa}$ 
such that $X_{\alphaa}$ is smooth at the generic point of $G$.
Let $v_G$ be the corresponding discrete valuation of $\CC(X)$.
Write 
\[
E_{\alphaa}\mid_{ X_{\alphaa}}=m_G G+(\text{other components}).
\]
Assume that $d_G=1$ and $G$ is not a toric subvariety 
in $\CC^n_{\alphaa}$. Then the discrepancy of $G$ with respect to $K_X$ 
is computed by the formula
\[
a(G,K_X)=|\alphaa|-1-v_{\alphaa}(\phi), \quad |\alphaa|=\sum \alpha_i,
\]
see \cite{Markushevich-1996-discr}. Let $S\subset X$ be a Cartier divisor and let $\psi$ 
be a local defining equation of $S$ in $\OOO_{0,X}$. 
Then $v_G(\psi)=v_{\alphaa}(\psi)$ and
the discrepancy of $G$ with respect to $K_X+cS$ 
is computed by the formula
\[
a(G,K_X+cS)=a(G,K_X)-cv_{G}(\psi)=|\alphaa|-1-v_{\alphaa}(\phi)-cv_{\alphaa}(\psi).
\]
Therefore,
\[
c \le a(G,K_X)/v_{\alphaa}(\psi)= (|\alphaa|-1-v_{\alphaa}(\phi))/v_{\alphaa}(\psi).
\]
\end{case}

\begin{definition}[cf. {\cite{Markushevich-1996-discr}}]
A weight $\alphaa$ is said to be \textit{admissible} if
$E_{\alphaa}\cap X_{\alphaa}$ contains at least one 
reduced non-toric component.
\end{definition}

\section{Gorenstein case}
In this section we consider the case where $(X\ni P)$ is either smooth or 
an index one singularity.
\begin{lemma}
\label{l-G-1}
If $(X\ni P)$ is smooth, then $c\le 5/6$.
\end{lemma}
\begin{proof}
Let $c>5/6$.
We may assume that $X=\CC^3$.
Let $\psi(x,y,z)=0$ be an equation of $S$.
Consider a weighted blowup 
$\sigma_{\alphaa}\colon \CC^3_{\alphaa}\to \CC^3$ with a suitable 
weight $\alphaa$. Let $E_\alphaa$ be the exceptional divisor.
Recall that $(S\ni P)$ is not Du Val.
Up to analytic coordinate change there are the following cases
(cf. \cite[4.25]{Kollar-Mori-19988}):

\begin{case} 
\textbf{Case $\ord_0 \psi\ge 3$.}
Take $\alphaa=(1,1,1)$ (usual blowup of $0$).
Then $a(E_{\alphaa}, K_X) =2$, $v_{\alphaa}(\psi)=\ord_0 \psi\ge 3$. Hence 
$c\le a(E_{\alphaa}, K_X)/v_{\alphaa}(\psi) \le 2/3$, a contradiction.
\end{case}

\begin{case} \textbf{Case $\psi=x^2+\eta(y,z)$, where $\ord_0 \eta\ge 4$.}
Take $\alphaa=(2,1,1)$.
Then $a(E_{\alphaa}, K_X) =3$, $v_{\alphaa}(\psi)=4$. Hence 
$c\le a(E_{\alphaa}, K_X) /v_{\alphaa}(\psi)\le 3/4$, a contradiction.
\end{case}

\begin{case} \textbf{Case $\psi=x^2+y^3+\eta(y,z)$, where $\ord_0 \eta\ge 4$.}
Here $\eta$ contains no terms $yz^l$, $l\le 3$ and $z^l$, $l\le 5$
(see, e.g., \cite[4.25]{Kollar-Mori-19988}).
Take $\alphaa=(3,2,1)$.
Then $a(E_{\alphaa}, K_X) =5$, $v_{\alphaa}(\psi)=6$. Hence 
$c\le a(E_{\alphaa}, K_X) /v_{\alphaa}(\psi)= 5/6$, a contradiction.
\end{case}
\end{proof}

\begin{lemma}
\label{l-G-2}
Assume that $(X\ni P)$ is a Gorenstein terminal singularity
and $(X\ni P)$ is not smooth. Then $c\le 4/5$.
\end{lemma}

\begin{proof}
Let $c>4/5$. 
We may assume that $X$ is a hypersurface in $\CC^4$
(it is an isolated cDV-singularity \cite{Reid-1980can}).
Let $\phi(x,y,z,t)=0$ be the equation of $X$.
Since $(X\ni P)$ is a cDV-singularity, 
$\ord_0 \phi=2$. According to \cite{Markushevich-1996-discr},
in a suitable coordinate system $(x,y,z,t)$, there is an admissible weighted blowup
$\sigma_{\alphaa} \colon \CC^4_{\alphaa}\to \CC^4$ such that at least 
for one component $G$ of $E_{\alphaa}\cap X_{\alphaa}$ 
we have $a(G,K_X)=1$. Then $c\le 1/v_{\alphaa}(\psi)$, so $v_{\alphaa}(\psi)=1$.
This means, in particular, that $\ord_0 \psi=1$. 
Up to coordinate change we may assume that
$\psi=t$. Write
\[
\phi=\eta(x,y,z)+t \zeta (x,y,z,t).
\]
Then $S$ is a hypersurface in $\CC^3_{x,y,z}$
given by $\eta(x,y,z)=0$.
As in the proof of Lemma \ref{l-G-1}, 
using Morse Lemma we get the following cases:

\begin{case} \textbf{Case $\ord_0 \eta\ge 3$.}
Take $\alphaa=(1,1,1,2)$. 
By the terminality condition \cite[Th. 4.6]{Reid-YPG1987}, 
we have $4=v_{\alphaa}(xyzt)-1>v_{\alphaa}(\phi)$.
Hence, $v_{\alphaa}(\eta)\le 3$ and $\eta_3\neq 0$.
We claim that $\alphaa$ is 
admissible whenever $\eta_3$ is not a cube of a linear form.
Indeed, in the affine chart $U_x:=\{x\neq 0\}$
the map $\sigma_{\alphaa}^{-1}$ is given by 
\begin{equation}
\label{eq-odd3}
x \mapsto x',\quad y \mapsto y'x',\quad z\mapsto z'x',\quad 
t \mapsto t'x'^2.
\end{equation}
First we assume that $\zeta$ contains the term $x$.
After the coordinate change $x\mapsfrom \zeta (x,y,z,t)$ we obtain
\[
\phi=\eta(x,y,z)+tx.
\]
Using \eqref{eq-odd3} we see that
$E_{\alphaa}\cap X_{\alphaa}$ is given in $\sigma_{\alphaa}^{-1}(U_x)\simeq 
\CC^4$ by 
\[
x'=\eta_3(1,y',z')+t'=0.
\]
Hence $\alphaa$ is admissible, i.e., 
$E_{\alphaa}\cap X_{\alphaa}$ has
a reduced non-toric component $G$. Then 
$a(G, K_X) =1$, 
$v_{G}(\psi)=2$ and $c\le a(G, K_X) /v_{G}(\psi)=1/2$, a contradiction.

Thus by symmetry we may assume that $\zeta$ contains no terms $x$, $y$, $z$.
Since $\ord_0 \phi=2$, $\zeta$ contains $t$. So, 
\[
\phi=\eta(x,y,z)+t^2+ t\xi (x,y,z,t),\quad \ord_0 \xi\ge 2.
\]
As above, $E_{\alphaa}\cap X_{\alphaa}$ is given in $\CC^4$ by 
$x'=\eta_3(1,y',z')=0$.
If $\eta_3$ is not a cube of a linear 
form, then $E_{\alphaa}\cap X_{\alphaa}$ has a reduced non-toric component $G$. 
Then, as above, 
$c\le 1/2$, a contradiction.

Consider the case where $\eta_3$ is a cube of a linear form. 
Then we may assume that 
$\eta_3(x,y,z)=y^3$, so 
\[
\phi=y^3+\eta^{\bullet}(x,y,z)+t^2+ t\xi (x,y,z,t),\quad \ord_0 \xi\ge 2, \quad 
\ord_0\eta^{\bullet} \ge 4.
\]
Put $\alphaa'=(2,2,2,3)$. Again, 
in the affine chart $U_x:=\{x\neq 0\}$
the map $\sigma_{\alphaa'}^{-1}$ is given by 
$x \mapsto x'^2$, $y \mapsto y'x'^2$, $z\mapsto z'x'^2$,
$t \mapsto t'x'^3$, 
where $\sigma_{\alphaa'}^{-1}(U_x)\simeq \CC^4/\muu_2(1,0,0,1)$ and 
\[
E_{\alphaa'}\cap X_{\alphaa'}\cap \sigma_{\alphaa'}^{-1}(U_x)=
\{ 
x'=0,\ y'^3+t'^2 =0\}.
\]
Thus $\alphaa'$ is admissible and for some component $G'$ of 
$X_{\alphaa'}\cap E_{\alphaa'}$ we have
$a(G', K_X)=2$, $v_{G'}(\psi)=3$, $c\le 2/3$, a contradiction.
\end{case}

\begin{case} \textbf{Case $\eta=x^2+\xi(y,z)$, where $\ord_0 \xi\ge 4$.}
By Morse Lemma we may assume that $\zeta$ does not depend on $x$.
Write $\zeta_1=\delta_1y+\delta_2z+\delta_3t$, $\delta_i\in \CC$.
Take $\alphaa=(2,1,1,3)$. 
In the affine chart $U_y:=\{y\neq 0\}$
the map $\sigma_{\alphaa}^{-1}$ is given by 
$x \mapsto x'y'^2$, $y \mapsto y'$, $z\mapsto z'y'$,
$t \mapsto t'y'^3$ and  
\[
E_{\alphaa}\cap X_{\alphaa}\cap \sigma_{\alphaa}^{-1}(U_y)=
\{ y'=0,\ x'^2+\xi_4(1,z')+\delta_1t' +\delta_2t'z'=0\}.
\]
If either $\delta_1\neq 0$ or $\delta_2\neq 0$ or $\xi_4\neq 0$, then 
$E_{\alphaa}\cap X_{\alphaa}$ is reduced (at least over $U_y$). 
Hence, 
$\alphaa$ is 
admissible and for some component $G$ of $E_{\alphaa}\cap X_{\alphaa}$
we have
$c\le a(G, K_X) /v_{G}(\psi)= 2/3$, a contradiction.
Thus $\delta_1=\delta_2= 0$ and $\xi_4=0$. 
Then we can write
\[
\phi=x^2+\xi(y,z)+\delta_3t^2+t\zeta^{\bullet}(y,z,t), \quad \ord_0\xi\ge 5,
\quad \ord_0\zeta^{\bullet}\ge 2.
\]
Take $\alphaa'=(2,1,1,2)$. 
In the affine chart $U_y:=\{y\neq 0\}$
the map $\sigma_{\alphaa'}^{-1}$ is given by 
$x \mapsto x'y'^2$, $y \mapsto y'$, $z\mapsto z'y'$,
$t \mapsto t'y'^2$ and  
\[
E_{\alphaa'}\cap X_{\alphaa'}\cap \sigma_{\alphaa'}^{-1}(U_y)=
\{ y'=0,\ x'^2+\delta_3t'^2+t\lambda(1,z')=0\},
\]
where $\lambda$ is the degree $2$ 
homogeneous part of $\zeta(y,z,0)$.
If $\delta_3\neq 0$ or $\lambda\neq 0$, 
as above, $\alphaa'$ is 
admissible and $c\le 1/2$, a contradiction.
Thus $\delta_3=0$, $\lambda= 0$, and
\[
\phi=x^2+\xi(y,z)+\delta t^3+t\zeta^{\circ}(y,z,t), \quad \delta\in \CC, \quad\ord_0\xi\ge 5,
\quad \ord_0\zeta^{\circ}\ge 3.
\]
Applying the terminality condition \cite[Th. 4.6]{Reid-YPG1987}
with weight $(2,1,1,1)$ we get that $\delta\neq 0$.

Take $\alphaa''= (3,1,1,2)$. As above we get that $\alphaa''$ 
is admissible and then
$c\le 1/2$, a contradiction.
\end{case}

\begin{case} \textbf{Case $\eta=x^2+y^3+\xi(y,z)$, where $\ord_0 \xi\ge 4$.}
Here $\xi$ contains no terms $yz^l$, $l\le 3$ and $z^l$, $l\le 5$
(see, e.g., \cite[4.25]{Kollar-Mori-19988}).
Write $\zeta_1=cz+\ell(x,y,t)$ and 
$\xi=\xi_{(6)}+\xi_{(7)}+\cdots$, where $\xi_{(d)}$ is 
the degree $d$ 
weighted homogeneous part of $\xi$ with respect to $\wt(y,z)=(2,1)$.
Here $\xi_{(6)}$ is a linear combination of $z^6$, $yz^4$,  $y^2z^2$.
Take $\alphaa=(3,2,1,5)$.
In the affine chart $U_z:=\{z\neq 0\}$
the map $\sigma_{\alphaa}^{-1}$ is given by 
$x \mapsto x'z'^3$, $y \mapsto y'z'^2$, $z\mapsto z'$,
$t \mapsto t'z'^5$ and
\[
E_{\alphaa}\cap X_{\alphaa}\cap \sigma_{\alphaa}^{-1}(U_z)=
\{z'=0,\ x'^2+y'^3+\xi_{(6)}(y',1)+\delta t'=0\},
\]
where $\delta$ is a constant and $\xi_{(6)}(y',1)$ contains no $y'^3$.
Hence $\alphaa$ is admissible, i.e., 
$E_{\alphaa}\cap X_{\alphaa}$ has
a reduced non-toric component $G$. Then 
$a(G, K_X) =4$, $v_{G}(\psi)=5$, and
$c\le a(G, K_X) /v_{G}(\psi)\le 4/5$, a contradiction.
\end{case}
\end{proof}

The following examples show that 
bounds $\ct(X,S)\le 5/6$ and $\le 4/5$ in 
Theorem \ref{main-2} are sharp.

\begin{example}
Let $X=\CC^3$ and let $S=S^d$ is given by $x^2+y^3+z^d$, $d\ge 6$.
Then $\ct(\CC^3,S^d)=5/6$. We prove this by descending induction on $\down{d/6}$.
Take $\alphaa=(3,2,1)$ and consider the $\alphaa$-blowup 
$\sigma_{\alphaa}\colon \CC^3_{\alphaa}\to \CC^3$.
Let $S_{\alphaa}\subset X_{\alphaa}$ be the proper transform of $S$.
We have $a(E_{\alphaa}, K_X) =5$ and $v_{\alphaa}(\psi)=6$. Hence, $\ct(\CC^3,S^d)\le 5/6$.
Further,
\[
\textstyle
K_{\CC^3_{\alphaa}}+\frac56S_{\alphaa}=\sigma_{\alphaa}^*(K_{\CC^3}+\frac56S).
\]
Thus it is sufficient to show that $\ct(X_{\alphaa},\frac 56 S_{\alphaa})$ is canonical.
We have three affine charts: 
\begin{itemize}
 \item 
$U_x:=\{x\neq 0\}$. Here $\sigma_{\alphaa}^{-1}\colon$ 
$x \mapsto x'^3$, $y \mapsto y'x'^2$, $z\mapsto z'x'$,
$S_{\alphaa}$ is given in $\sigma_{\alphaa}^{-1}(U_x)\simeq\CC^3/\muu_3(-1,2,1)$ by the equation 
$1+y'^3+z'^dx'^{d-6}=0$. Hence, in this chart, 
$S_{\alphaa}$ is smooth and does not pass through a (unique) singular point
of $\sigma_{\alphaa}^{-1}(U_x)$.
 \item 
$U_y:=\{y\neq 0\}$. Here $\sigma_{\alphaa}^{-1}\colon$ 
$x \mapsto x'y'^3$, $y \mapsto y'^2$, $z\mapsto z'y'$,
$S_{\alphaa}$ is given in $\sigma_{\alphaa}^{-1}(U_y)\simeq \CC^3/\muu_2(3,-1,1)$ by the equation 
$x'^2+1+z'^dy'^{d-6}=0$. Again, in this chart, 
$S_{\alphaa}$ is smooth and does not pass through a (unique) singular point
of $\sigma_{\alphaa}^{-1}(U_y)$.

 \item 
$U_z:=\{z\neq 0\}$. Here $\sigma_{\alphaa}^{-1}\colon$ 
$x \mapsto x'z'^3$, $y \mapsto y'z'^2$, $z\mapsto z'$,
$S_{\alphaa}$ is given in $\sigma_{\alphaa}^{-1}(U_z)\simeq \CC^3$ by the equation 
$x'^2+y'^3+z'^{d-6}=0$. In this chart, $(X_{\alphaa},S_{\alphaa})\simeq (\CC^3, S^{d-6})$.
\end{itemize}
Thus $X_{\alphaa}$ has only terminal singularities,
$S_{\alphaa}$ does not pass through any singular point of 
$X_{\alphaa}$, 
and the pair $(X_{\alphaa}, S_{\alphaa})$ is terminal in
charts $U_x$ and $U_y$. In the chart $U_z$ the pair
by induction $(X_{\alphaa},\frac 56 S_{\alphaa})$ is canonical
(moreover, $(X_{\alphaa}, S_{\alphaa})$ is canonical if $d\le 11$).
Therefore, $\ct(X,S)=5/6$.
\end{example}

\begin{example}
Let $X\subset \CC^4$ is given by 
$x^2+y^3+z^d+tz=0$, $d\ge 7$ and let $S$ cut out by $t=0$.
Take $\alphaa=(3,2,1,5)$ and
consider the $\alphaa$-blowup $\sigma_{\alphaa} \colon X_{\alphaa}\to X$.
Let $S_{\alphaa}\subset X_{\alphaa}$ be the proper transform of $S$.
We see below that $\alphaa$ is admissible.
Moreover, the exceptional divisor $G:=E_{\alphaa}\cap X_{\alphaa}$
is reduced and irreducible.
We have four charts: 
\begin{itemize}
 \item 
$U_x:=\{x\neq 0\}$. Here $\sigma_{\alphaa}^{-1}\colon
x \mapsto x^3$, $y \mapsto yx^2$, $z\mapsto zx$, $t\mapsto tx^5$,
$X_{\alphaa}$ is given in $\sigma_{\alphaa}^{-1}(U_x)\simeq\CC^4/\muu_3(-1,2,1,5)$ by the equation 
$1+y^3+z^dx^{d-6}+tz=0$ and $S_{\alphaa}$ by two equations $x=1+y^3+tz=0$. 
Hence, in this chart, both $X_{\alphaa}$ and
$S_{\alphaa}$ are smooth. 
 \item 
$U_y:=\{y\neq 0\}$. Here
$\sigma_{\alphaa}^{-1}\colon 
x \mapsto xy^3$, $y \mapsto y^2$, $z\mapsto zy$, $t\mapsto ty^5$,
$\sigma_{\alphaa}^{-1}(U_y)\simeq \CC^4/\muu_2(3,-1,1,5)$, 
$X_{\alphaa}=\{x^2+1+z^dy^{d-6}+tz=0\}$, and $S_{\alphaa}=\{y=x^2+1+tz=0\}$. 
As above, both $X_{\alphaa}$ and
$S_{\alphaa}$ are smooth in this chart.
 \item 
$U_z:=\{z\neq 0\}$. Here $\sigma_{\alphaa}^{-1}\colon
x \mapsto xz^3$, $y \mapsto yz^2$, $z\mapsto z$, $t\mapsto tz^5$,
$\sigma_{\alphaa}^{-1}(U_z)\simeq \CC^4$,   
$X_{\alphaa}=\{x^2+y^3+z^{d-6}+t=0\}$, and $S_{\alphaa}=\{z=x^2+y^3+t=0\}$. 
As above, both $X_{\alphaa}$ and
$S_{\alphaa}$ are smooth in this chart.

 \item 
$U_t:=\{t\neq 0\}$. Here $\sigma_{\alphaa}^{-1}\colon
x \mapsto xt^3$, $y \mapsto yt^2$, $z\mapsto zt$, $t\mapsto t^5$,
$\sigma_{\alphaa}^{-1}(U_t)\simeq \CC^4/\muu_5(3,2,1,-1)$,
$X_{\alphaa}=\{x^2+y^3+z^dt^{d-6}+z=0\}$, and $S_{\alphaa}=\{t=x^2+y^3+z=0\}$. 
The variety
$X_{\alphaa}$ has a unique singular point $Q$ at the origin and this point is terminal of type 
$\frac15(3,2,-1)$ the surface
$S_{\alphaa}$ is smooth and does not pass through $Q$. 
\end{itemize}
Thus we have $a(G, K_X) =4$, $v_{\alphaa}(\psi)=5$, and $a(G, K_X+ \frac45 S) =0$.
Therefore, 
\[
\textstyle
K_{X_{\alphaa}}+\frac45S_{\alphaa}=\sigma_{\alphaa}^*(K_{X}+\frac45S).
\]
Since the pair $K_{X_{\alphaa}}+\frac45S_{\alphaa}$ is canonical, $\ct(X,S)=4/5$.
\end{example}

\section{Non-Gorenstein case}
Now we assume that $(X\ni P)$ is a (terminal) point of index $r>1$.
Let $\pi \colon (X^\sharp\ni P^\sharp)\to (X\ni P)$ be the index-one
cover and let $S^\sharp:=\pi^{-1}(S)$.

\begin{lemma}
\label{l-nG-1}
If $(X\ni P)$ is a cyclic quotient singularity, then $\ct(X,S)\le 1/2$.
\end{lemma}

\begin{proof}
By our assumption we have
$X\simeq \CC^3/\muu_r(a,-a,1)$ for some $r\ge2$, $1\le a <r$, $\gcd(a,r)=1$.
Assume that $c=\ct(X,S)> 1/2$. 
Let 
$\psi=0$ be a defining equation of $S^\sharp$.
Consider the weighted blowup $\sigma_{\alphaa}\colon X_{\alphaa}\to X$
with weights $\alphaa =\frac1r (a,r-a,1)$.
Then $a(E_{\alphaa}, K_X)=1/r$.
Since $a(E_{\alphaa}, K_X)-c v_\alphaa (\psi)\ge 0$, we have 
$v_\alphaa (\psi)\le a(E_{\alphaa}, K_X)/c< 2a(E_{\alphaa}, K_X)= 2/r$
and so $v_\alphaa (\psi)=1/r$.
Thus we may assume that $\psi$ contains $x_3$
(if $a\equiv \pm 1$ we possibly have to permute coordinates).
Then $S^\sharp\simeq \CC^2$ is smooth and $S\simeq \CC^2/\muu_r(a,-a)$, i.e.,
$S$ is Du Val of type $A_{r-1}$.
\end{proof}

\begin{lemma}
\label{l-nG-2}
If $(X\ni P)$ is a terminal singularity of index $r>1$ 
and $\ct(X,S)>1/2$, then $K_X+S\sim 0$.
\end{lemma}

\begin{proof}
By Lemma \ref{l-nG-1} $(X\ni P)$ is not a cyclic quotient singularity.
There is an analytic $\muu_r$-equivariant 
embedding $(X^\sharp, P^\sharp) \subset (\CC^4,0)$. 
Let $(x_1,x_2,x_3,x_4)$ be coordinates in $\CC^4$, let $\phi=0$ be an 
equation of $X^\sharp$, and let $\psi=0$ be an equation of $S^{\sharp}$. 
We can take $(x_1,x_2,x_3,x_4)$ and $\phi$ 
to be semi-invariants 
such that one of the following holds \cite{Reid-YPG1987}:
\begin{enumerate}
\item[-] 
\textbf{Main series.}
$\wt (x_1,x_2,x_3,x_4; \phi)\equiv (a,-a, 1, 0;0)\mod r$, where $\gcd (a,r)=1$.
\item[-] 
\textbf{Case $cAx/4$.} $r=4$,
$\wt (x_1,x_2,x_3,x_4; \phi)\equiv (1,3, 1, 2;2)\mod 4$.
\end{enumerate}
In both cases $\wt (x_1x_2x_3x_4)- \wt \phi\equiv \wt x_3\mod r$.
According to \cite{Kawamata-1992-discr} there is a weight 
$\alphaa$
such that for the corresponding $\alphaa$-blowup
$\sigma_{\alphaa}\colon X_{\alphaa}\subset W\to X\subset \CC^4/\muu_r$ 
the exceptional divisor $E_{\alphaa}\cap X_{\alphaa}$ 
has a reduced component $G$ of 
discrepancy $a(G, K_X)=1/r$.
Moreover, $r\alpha_i\equiv \wt x_i \mod r$, $i=1,2,3,4$.
Since $c>1/2$, we have $1/r-cv_\alphaa (\psi)\ge 0$, i.e., 
$rv_\alphaa (\psi)< 2$, so $rv_\alphaa (\psi)=1$. 
In particular, $\wt \psi\equiv 1 \mod r$. 

Let $\omega$ 
be a section of $\OOO_X(-K_X)$. Then $\omega$ can be written as 
\[
\omega=\lambda (\partial \phi/\partial x_4)(dx_1 \wedge dx_2\wedge dx_3)^{-1},
\]
where
$\lambda$ is a semi-invariant function with 
\[
\wt \lambda-\wt (x_1x_2x_3x_4)+\wt \phi\equiv \wt\omega \equiv 0\mod r.
\]
Thus, $\wt \psi \equiv \wt \lambda\mod r$. Hence, $S\sim -K_X$.
\end{proof}

\begin{lemma}
\label{l-nG-3}
If $(X\ni P)$ is a terminal singularity of index $r>1$, 
then $c\le 4/5$.
\end{lemma}

\begin{proof}
Since $\pi$ is \'etale in codimension one, 
we have $K_{X^\sharp }+cS^\sharp=\pi^*(K_X+cS)$. Hence
the pair $(X^\sharp,\, cS^\sharp)$ is canonical (see, e.g., 
\cite[3.16.1]{Kollar-1995-pairs}). Assume that $c> 4/5$.
By Lemma \ref{l-nG-1} the point $(X^\sharp\ni P^\sharp)$ is singular.
Then by Lemma \ref{l-G-2} the pair 
$(X^\sharp,\, S^\sharp)$ is canonical.
Therefore, $(S^\sharp\ni P^\sharp)$ is a Du Val singularity.
Then the singularity $(S\ni P)=(S^\sharp\ni P^\sharp)/\muu_r$
is log terminal. On the other hand, by Lemma \ref{l-nG-2}
the divisor $K_S$ is Cartier. Hence, $(S\ni P)$ is Du Val, a contradiction.
\end{proof}

%\bibliography{my_ref}
%\bibliographystyle{alpha}%{plain}%{alpha}%%{unrst} {abbrv}

\end{document}